How many gallons of a 4% salt solution must be mixed with 25 gallons of a 6% solution to obtain a 5% solution?
Answer by richwmiller(17219)

0.04*x+0.06*25=0.05(25+x)
0.04*x+1.5=1.25+0.05x
0.04x-0.05x=1.25-1.5
-0.01x=-0.25
x=25 gallons of 4% is added to 25 gallons of 6% to get 50 gallons at 5%
check
0.04*25+0.06*25=0.05(25+25)
1+1.5=0.05(50)
2.5=2.5
ok